OpenStudy (anonymous):

During a rescue operation, a 5300kg helicopter hovers at a fixed height. The rotor blades send air downward at a speed of 62m/s. What mass of air must pass through the blades every second to produce enough thrust to hold the helicopter in place?

OpenStudy (anonymous):

 Sistem: Helicopter(mass: 5300 Kg) Surroundings(Earth, Air) Initial State: vi=62 m/s Final State: FBD:  F = mass * acerelation Fgrav = mHelicopter * g Fgrav = (5300 Kg)(-9.8 m/s^2) Fgrav = -51.94 m/s^2 Fblades(helicopter) = mAir * aAir Fblases/vAir = mAir Fgrav = Fblades(helicopter) Proportial (ie: equilibrium. hovering) mAir = 51,94 m/s^2 / 62 m/s = 873.74 Kg/s^2 Therefore 873.74 Kg of air pass through the blades every second for the helicopter to remain in the air (hovering).

OpenStudy (anonymous):

I made drawings and the FBD but OpenStudy lag so much that didnt let me post it, so I had to refresh and copy the whole solution again... Still it works.